### v1

# Electron polarization induced by electroweak interaction in high-energy scattering off light nuclei

Abstract
Electron polarization plays a significant role in studies on nuclear scattering. Nevertheless, the development of a comprehensive approach to such a problem remains challenging, particularly at the relativistic electron-energy scale. Herein, we present a theoretical approach to investigate the impact of electron polarization in scattering off unoriented light nuclei, based on the multipole expansion for the scattering cross section at high energies. Numerical calculations for stable 6,7Li and unstable 7Be nuclei show that the longitudinal polarization and weak interaction are not explicitly correlated when electrons scatter at 0 for all energy scales. In contrast, their correlation is strongly exhibited at other scattering angles when energy exceeds 10 GeV. A comparison between stable and unstable nuclei reveals that the stability of nuclei significantly affects the electron polarization contribution to the scattering cross section. This study opens new approaches to the nuclear structure problems, in particular the EMC effect, via the deep inelastic electron-nucleus scattering using the unified electroweak theory.
Technical advancements in generating polarized electron beams enable us to test the predictions from theoretical models and obtain additional data regarding their scattering on both oriented and unoriented nuclei [1, 2, 3, 4, 5, 6, 7, 8]. Experimentally, the polarization of an electron beam can either remain unchanged or be altered after scattering with a target [2, 5]. This means that the scattering cross section and asymmetry depend on both the magnitude of electron polarization and how it changes [2, 3, 6, 8]. One must either calculate the polarization density matrix through the cross section or precisely produce the polarized beams for experimental use [1, 4, 9, 10]. Scattering processes at high energies are often of particular interest since electrons have only longitudinal polarization in the direction of momentum transfer [2].
It was initiated long ago that the multipole expansion for the electromagnetic cross section can be applied to examine the asymmetry resulting from electron polarization and nuclear spin orientation, and clarify their impact [1, 2]. A similar expansion for the electroweak one, which includes the contribution of weak interaction, has also been proposed [11, 12, 13, 14]. The nuclear form factors in the expressions in prior studies [1, 2, 11, 12] were either approximated or treated as parameters, while they were precisely calculated within later ones [13, 14] using the unified electroweak theory combined with the many-particle shell model and fractional parentage coefficients [15, 16, 17]. In particular, by employing the V-A structure of the neutral weak current along with appropriately chosen harmonic oscillator parameters, the latest study in Ref. [14] has successfully described the experimental data for the unpolarized electron scattering on 6,7Li and 7Be nuclei at MeV energies [18, 19], and provided predictive insights at GeV energies.
To date, the study of high-energy electron polarization variability and its profound impact on electroweak scattering processes, which reveal information beyond what is inaccessible through pure electromagnetic interaction or unpolarized electron sources, has not been thoroughly investigated. Such a study also points towards a novel approach, based on the unified electroweak theory, for unveiling the origin of the EMC effect discovered in deep inelastic scattering (DIS) experiments [20, 21, 22]. The present Letter presents the first comprehensive investigation of all possible longitudinally polarized electron scatterings from unoriented light stable and unstable nuclei using the multipole expansion technique within the framework of the unified electroweak theory. We also examine the impact of electron polarization and its correlation with the weak interaction. This provides a basis for the future studies on asymmetry resulting solely from electron polarization and the weak interaction, once nuclear spin orientation is excluded.
The total electron-nucleus scattering cross section in the multipole expansion within the unified electroweak theory can be written as [13, 14]:
$\sigma_{\text{total}}^{\text{pol.}}(\xi,\xi^{\prime})=\sigma_{\text{total}}^{0}+\sigma_{\text{total}}^{\xi\xi^{\prime}},\vspace{-20pt}$ (1)
where $\xi$ and $\xi^{\prime}$ are the electron polarizations before and after scattering, whose values are equal to +1 or $-1$. The first term in Eq. (1), which does not involve electron polarization, is a sum of
$\sigma_{F}^{0}$ $=\frac{\alpha^{2}\varepsilon^{\prime}}{f\varepsilon Q^{4}}A_{1},$ (2a)
$\sigma_{FZ}^{0}$ $=\frac{2\alpha\lambda\varepsilon^{\prime}}{f\varepsilon Q^{4}}(g_{V}B_{1}+g_{A}B_{2}),$ (2b)
$\sigma_{Z}^{0}$ $=\frac{\lambda^{2}\varepsilon^{\prime}}{f\varepsilon Q^{4}}\left[\left(g_{V}^{2}+g_{A}^{2}\right)C_{1}+2g_{V}g_{A}C_{2}\right],$ (2c)
while the second term is a respective sum of
$\sigma_{F}^{\xi\xi^{\prime}}$ $=\frac{\alpha^{2}\varepsilon^{\prime}}{f\varepsilon Q^{4}}[\xi\xi^{\prime}A_{1}+(\xi+\xi^{\prime})A_{2}],$ (3a)
$\sigma_{FZ}^{\xi\xi^{\prime}}$ $=\frac{2\alpha\lambda\varepsilon^{\prime}}{f\varepsilon Q^{4}}\{[g_{V}\xi\xi^{\prime}+g_{A}(\xi+\xi^{\prime})]B_{1}$
$\phantom{000000.}+[g_{V}(\xi+\xi^{\prime})+g_{A}\xi\xi^{\prime}]B_{2}\},$ (3b)
$\sigma_{Z}^{\xi\xi^{\prime}}$ $=\frac{\lambda^{2}\varepsilon^{\prime}}{f\varepsilon Q^{4}}\{[(g_{V}^{2}+g_{A}^{2})\xi\xi^{\prime}+2g_{V}g_{A}(\xi+\xi^{\prime})]C_{1}$
$\phantom{000000}+[(g_{V}^{2}+g_{A}^{2})(\xi+\xi^{\prime})+2g_{V}g_{A}\xi\xi^{\prime}]C_{2}\}.$ (3c)
In Eqs. (2a)–(3c), $\varepsilon$ and $\varepsilon^{\prime}$ are the electron energies before and after scattering, respectively; $Q=K-K^{\prime}=(\omega,\mathbf{q})$ is the momentum transfer with $K=(\varepsilon,\mathbf{k})$ and $K^{\prime}=(\varepsilon^{\prime},\mathbf{k^{\prime}})$ being the electron momenta. $\alpha$ and $g$ are the electromagnetic and weak coupling constants, and $f$ is the nuclear recoil factor. $\lambda=-G_{F}m_{Z}^{2}Q^{2}/[2\sqrt{2}\pi(m_{Z}^{2}-Q^{2})],G_{F}=g^{2}/4\sqrt{2}m_{Z}^{2}\cos^{2}\theta_{\text{W}}$ with $\theta_{\text{W}}$ being the Weinberg angle, and $m_{Z}$ being the $Z^{0}$ boson mass. $g_{V}=-1/2+2x_{\text{W}},g_{A}=-1/2$ and $x_{\text{W}}\equiv\sin^{2}\theta_{\text{W}}$ are the weakinteraction parameters in the Weinberg-Salam model. The coefficients $A_{1},A_{2},B_{1},B_{2},C_{1}$, and $C_{2}$ are calculated in terms of the multipole form factors, whose explicit forms can be found in Ref. [13].
Obviously, the polarized cross sections [Eqs. (3a)–(3c)] depend on the polarization degrees $\xi$ and $\xi^{\prime}$. For simplicity, we assign $\xi\,(\xi^{\prime})=+1$ for the right-handed polarization and $\xi\,(\xi^{\prime})=-1$ for the left-handed one. We then consider all possible variations of electron polarization as following cases. Case 1: both incoming and outgoing electron beams are unpolarized. This case has been studied in detail in our recent study [14]. Case 2: electron polarizations before and after scattering are preserved, i.e., $\xi=\xi^{\prime}$. Case 3: an initially polarized beam becomes unpolarized, i.e., $\xi=+1/-1$ and $\xi^{\prime}=0$ (or $\xi^{\prime}$ is taken over +1 and $-1$). Case 4: an electron beam reverses its initial polarization after scattering, i.e., $\xi=-\xi^{\prime}$.
The total cross section (1) will be equivalent to that of unpolarized electron scattering when taking the summation over all possible electron polarization states. This indicates that an unpolarized electron beam has the spin components uniformly distributed in all directions [5]. We can also see that, if $\xi=\xi^{\prime}=+1/-1$, the electron polarization contribution via the pure electromagnetic term [Eq. (3a)] to the total cross section is identical to the unpolarized term [Eq. (2a)], namely, $\sigma_{F}^{11}=\sigma_{F}^{-1-1}=\sigma_{F}^{0}$. This proves that electron polarization is preserved under the electromagnetic interaction, as previously mentioned in Ref. [2]. In contrast, the contributions via the terms involving the weak interaction [Eqs. (3b)–(3c)] to the total cross section are unequal and different from the respective unpolarized ones [Eqs. (2b)–(2c)]. Particularly, as $\xi=+1/-1$ and $\xi^{\prime}=0$, only the last two terms (3b) and (3c) make a nonzero contribution to the total cross section. This reveals that electron polarization can be changed only when the weak interaction is present, while it is unchanged under the electromagnetic one. It is also easy to find a similar result for the inverse case, i.e., $\sigma_{F}^{0\pm 1}=\sigma_{F}^{\pm 10}=0$, while $\sigma_{FZ}^{0\pm 1}=\sigma_{FZ}^{\pm 10}\neq 0$ and $\sigma_{Z}^{0\pm 1}=\sigma_{Z}^{\pm 10}\neq 0$. So, we conclude that the weak interaction can make an electron beam with the $\xi=+1/-1$ initial polarization become an unpolarized beam after scattering, and vice versa. However, the total cross section (1) will vanish when the electron polarizations before and after scattering have opposite signs. This implies that the electroweak interaction cannot cause the spin orientation of all electrons to be reversed, but alters half of it. The specific expressions corresponding to all cases under discussion can be readily derived from Eqs. (3a)–(3c).
Now, we examine two typical polarized scattering cross sections, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ and $\sigma_{\text{total}}^{\text{pol.}}(1,1)$, and compare them to the unpolarized one $\sigma_{\text{total}}^{\text{unpol.}}$, where $\sigma_{\text{total}}^{\text{unpol.}}=2\sigma_{\text{total}}^{0}$ as aforementioned [13, 14, 17]. Here, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ reveals that the incoming right-handed polarization electron beam becomes unpolarized after scattering. Similarly, $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ implies both incoming and outgoing electron beams being the right-handed polarization ones. Using multipole expansion for the scattering cross section, such calculations can be easily carried out once all nuclear form factors are well defined. For practical calculations, we investigate the elastic scattering of polarized electrons on ${}^{6,7}\text{Li}$ and ${}^{7}\text{Be}$ nuclei, whose form factors and other relevant parameters are taken from Ref. [14]. Numerical calculations are performed with various incident electron energies and scattering angles.
Fig. 1 shows the calculated $\sigma_{\text{total}}^{\text{pol.}}(1,0)/\sigma_{\text{total}}^{\text{unpol.}}$ ratio for 6Li, whereas Fig. 2 plots the computed $\sigma_{\text{total}}^{\text{pol.}}(1,1)/\sigma_{\text{total}}^{\text{unpol.}}$ one for 7Li. The $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ ratio for 7Be is also displayed in Fig. 3. All the crosssection ratios for three considered nuclei are listed in Tab. I. It can be seen that at energies not exceeding $10$ GeV, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ is just half of $\sigma_{\text{total}}^{\text{unpol.}}$, while $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is almost equal to the latter. A slight difference is observed when comparing the cross-section ratios for three nuclei (Tab. I). This can be attributed to the negligible contribution of weak interaction to the polarized terms at such an energy range [13]. When the electron energy goes above 10 MeV, the weak interaction gradually emerges and correlates with electron polarization to make polarized cross sections considerably increase, as seen in Figs. 1–3.
Figure 1: The polarized cross sections $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ versus the unpolarized ones $\sigma_{\text{total}}^{\text{unpol.}}$ obtained for 6Li at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and scattering angles $\theta$.
We note that the exponential functions in the numerator and denominator cancel out each other when calculating all considered ratios, while the scattering cross section itself rapidly decreases toward zero with increasing scattering angles at high energies, as previously reported in Refs. [13, 14]. Therefore, their plots remain fully visible across all scattering angles. However, our data table only lists the ratios at narrow angles ($\theta=0^{\circ}-8^{\circ}$) since their values at larger angles are not particularly meaningful once all the scattering cross sections are extremely small. The results in Tab. I also indicate that the information obtained in each scattering process strongly depends on how electron polarization changes. Specifically, once electron polarization remains unchanged, the scattering cross section ($\sigma_{\text{total}}^{\text{pol.}}(1,1)$) is significantly larger than that obtained when the outgoing beam becomes unpolarized ($\sigma_{\text{total}}^{\text{pol.}}(1,0)$), in particular at energies above 100 GeV and scattering angles $\theta=5^{\circ}-8^{\circ}$.
Figure 2: The polarized cross sections $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ versus the unpolarized ones $\sigma_{\text{total}}^{\text{unpol.}}$ obtained for 7Li at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and scattering angles $\theta$.
Since the unpolarized term $\sigma_{\text{total}}^{0}$ in the total polarized cross section is identical in all cases, we can make a direct comparison between $\sigma_{\text{total}}^{10}$ and $\sigma_{\text{total}}^{11}$, whose calculated ratios for three nuclei are also provided in Tab. I. At energies below 10 GeV, the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratios are very small, with the highest order of magnitude being only around $10^{-4}$ at $\theta=8^{\circ}$. However, at higher energy scales, these ratios increase remarkably, reaching up to several tens of percent. At the same energy and scattering angle, except for $\theta=0^{\circ}$, their magnitudes follow an ascending order of 7Li, 6Li, and 7Be nuclei.
To further clarify the impact of electron polarization in each scattering process, let us now consider the ratios $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ shown in Tab. I. Similar to the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratio, the contribution of $\sigma_{\text{total}}^{10}$ can be neglected at 10 GeV, since its highest order of magnitude relative to $\sigma_{\text{total}}^{0}$ at $\theta=8^{\circ}$ is only around $10^{-4}$. In contrast, $\sigma_{\text{total}}^{11}$ is nearly equal to $\sigma_{\text{total}}^{0}$ at energies not more than 10 GeV. At higher energy scales, the contribution of $\sigma_{\text{total}}^{10}$ becomes significant, whereas $\sigma_{\text{total}}^{11}$ even exceeds $\sigma_{\text{total}}^{0}$ by several times, as shown in Fig. 3 for 7Be.
Figure 3: The polarized components $\sigma_{\text{total}}^{11}$ versus the unpolarized ones $\sigma_{\text{total}}^{0}$ in the total polarized cross sections obtained for 7Be at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and scattering angles $\theta$.
Overall, all the above ratios show a little difference among selected nuclei in the scattering processes at energies below 10 GeV. At higher energies, their magnitudes increase in the order of 7Li, 6Li, and 7Be nuclei. Particularly, there are remarkable discrepancies in the contribution of electron polarization to the scattering cross section when comparing stable and unstable nuclei. The 7Li -7Be pair proves this remark, although they have the same nucleon numbers. In contrast, for the 6Li -7Li stable pair, these discrepancies are much smaller, even though their nucleon numbers are slightly different. It can be explained that the weak interaction between electrons and nuclei strongly depends on the nuclear stability. This property affects the correlation between the weak interaction and electron polarization, along with their contribution in each scattering process.
Furthermore, although the $\sigma_{\text{total}}^{10}$ term can take negative values, its magnitude is significantly smaller than the $\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}$ ones. Therefore, the total cross section also depends on the correlation between the electron spin orientations before and after scattering and remains positive in all cases. In particular, when the longitudinally polarized electrons move straight forward after scattering, the electron polarization-weak interaction correlation is not exhibited. This is evidenced by the ratios at $\theta=0^{\circ}$ in Tab. I. Their values remain nearly unchanged at all considered energies, except for $\sigma_{\text{total}}^{10}$, whose contribution can be omitted. So, besides nuclear spin orientation, it is evident that electron polarization and its transformation manner under the weak interaction play a crucial role in generating various types of asymmetry in electron-nucleus scattering at high energies.
In conclusion, we have performed a systematic study on the impact of electron polarization on scattering off unoriented light nuclei using the unified electroweak theory. Numerical calculations of the scattering cross sections in different electron polarization cases for three 6,7Li and 7Be light nuclei are obtained by directly calculating their multipole form factors, using the formulas provided in Ref. [14]. The results obtained are then compared and/or correlated with the corresponding cross sections for the unpolarized electron scattering case. It has been found that at the scattering angle of $\theta=0^{\circ}$, the longitudinal electron polarization does not explicitly correlate with the weak interaction for all energy scales. Such a correlation is strongly exhibited at nonzero scattering angles ($\theta=5^{\circ}-8^{\circ}$) and high energies above 10 GeV. By comparing the results obtained for stable and unstable nuclei, we found that the nuclear stability significantly influences the contribution of electron polarization to the scattering cross section. This research direction can be extended in many ways, specifically considering various state transitions within a single nucleus or different nuclei. Then, the asymmetry resulting solely from electron polarization and the weak interaction can be figured out. In addition, other interesting results are expected once the nuclear spin orientation and charge-changing weak current are included. Moreover, the multipole expansion technique for the scattering cross section within the unified electroweak theory also suggests a new approach to directly study the EMC effect in a target nucleus using both unpolarized and polarized electrons. More specifically, we calculate the accumulative cross section contributing by separate nucleons inside a certain nucleus and compare it with that made of a similar number of nucleons bound inside the same one. It means that we can investigate the above effect upon only a single nucleus instead of indirectly comparing its scattering cross section with that of a lighter nucleus. These extensive studies will provide further vital information, contributing to a thorough understanding on the nature of the EMC effect in deep inelastic electron-nucleus scattering.
Table 1: Calculated cross-section ratios for 6,7Li and 7Be obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and small scattering angles.
$\varepsilon$ $\theta$ $\sigma^{\text{pol.}}_{\text{total}}(1,0)/\sigma^{\text{unpol.}}_{\text{total}}$ $\sigma^{\text{pol.}}_{\text{total}}(1,1)/\sigma^{\text{unpol.}}_{\text{total}}$ $\sigma^{10}_{\text{total}}/\sigma^{11}_{\text{total}}$ $\sigma^{10}_{\text{total}}/\sigma^{0}_{\text{total}}$ $\sigma^{11}_{\text{total}}/\sigma^{0}_{\text{total}}$
(GeV) 7Li 6Li 7Be 7Li 6Li 7Be 7Li 6Li 7Be 7Li 6Li 7Be 7Li 6Li 7Be
10 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8237 ×$\times 10^{-14}$ 4.6965 ×$\times 10^{-12}$ 2.1329 ×$\times 10^{-12}$ 9.8237 ×$\times 10^{-14}$ 4.6965 ×$\times 10^{-12}$ 2.1329 ×$\times 10^{-12}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 4.2043 ×$\times 10^{-5}$ 3.2235 ×$\times 10^{-5}$ 8.4485 ×$\times 10^{-6}$ 4.2047 ×$\times 10^{-5}$ 3.2237 ×$\times 10^{-5}$ 8.4486 ×$\times 10^{-6}$ 1.0001 1.0001 1.0000
$5^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0001 1.$0$2.4903 ×$\times 10^{-5}$ 8.9254 ×$\times 10^{-5}$1.4668 ×$\times 10^{-5}$2.4902 ×$\times 10^{-5}$ 8.9270 ×$\times 10^{-5}$1.4667 ×$\times 10^{-5}$ 1.0000 1.0002 1.0000
$8^{\circ}$ 0.4999 0.5001 0.4998 0.9998 1.0002 0.9997 -2.0072 ×$\times 10^{-4}$ 2.2865 ×$\times 10^{-4}$3.4970 ×$\times 10^{-4}$2.0063 ×$\times 10^{-4}$ 2.2876 ×$\times 10^{-4}$3.4946 ×$\times 10^{-4}$ 0.9996 1.0005 0.9993
100 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8237 ×$\times 10^{-12}$ 4.6965 ×$\times 10^{-10}$ 2.1329 ×$\times 10^{-10}$ 9.8237 ×$\times 10^{-12}$ 4.6965 ×$\times 10^{-10}$ 2.1329 ×$\times 10^{-10}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5008 0.5016 0.5028 1.0016 1.0032 1.0055 0.0016 0.0032 0.0055 0.0016 0.0032 0.0055 1.0032 1.0064 1.0111
$5^{\circ}$ 0.5023 0.5044 0.5082 1.0045 1.0088 1.0165 0.0045 0.0087 0.0159 0.0045 0.0088 0.0165 1.0090 1.0177 1.0329
$8^{\circ}$ 0.5054 0.5112 0.5219 1.0109 1.0223 1.0438 0.0106 0.0214 0.0403 0.0109 0.0223 0.0438 1.0218 1.0447 1.0877
500 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 2.4559 ×$\times 10^{-10}$ 1.1741 ×$\times 10^{-8}$ 5.3315 ×$\times 10^{-9}$ 2.4559 ×$\times 10^{-10}$ 1.1741 ×$\times 10^{-8}$ 5.3315 ×$\times 10^{-9}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5173 0.5371 0.5753 1.0346 1.0741 1.1506 0.0324 0.0645 0.1157 0.0346 0.0741 0.1506 1.0692 1.1482 1.3011
$5^{\circ}$ 0.5414 0.5894 0.6793 1.0829 1.1788 1.3587 0.0711 0.1317 0.2088 0.0829 0.1788 0.3587 1.1657 1.3577 1.7173
$8^{\circ}$ 0.5811 0.6722 0.8217 1.1622 1.3444 1.6435 0.1225 0.2039 0.2814 0.1622 0.3444 0.6435 1.3244 1.6889 2.2870
1000 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8236 ×$\times 10^{-10}$ 4.6965 ×$\times 10^{-8}$ 2.1317 ×$\times 10^{-8}$ 9.8236 ×$\times 10^{-10}$ 4.6965 ×$\times 10^{-8}$ 2.1317 ×$\times 10^{-8}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5550 0.6183 0.7326 1.1100 1.2365 1.4652 0.0901 0.1606 0.2410 0.1100 0.2365 0.4652 1.2199 1.4730 1.9305
$5^{\circ}$ 0.6041 0.7173 0.8837 1.2081 1.4346 1.7674 0.1470 0.2325 0.3027 0.2081 0.4346 0.7674 1.4162 1.8693 2.5348
$8^{\circ}$ 0.6500 0.7995 0.9636 1.3000 1.5989 1.9272 0.1875 0.2725 0.3248 0.3000 0.5989 0.9272 1.6000 2.1979 2.8544
M. T. Vo would like to thank Dr. Sc. Luong Zuyen Phu for his useful discussions on this work.